OpenStudy (anonymous):

Multiply the two polynomials and write you answer in standard form. -3x(4x-2)

Nnesha (nnesha):

distribute parentheses by -3x for example \[\huge\rm \color{red}{a}(b+c) =\color{red}{a}b+\color{red}{a}c\] multiply everything which is inside the parentheses by outside term

OpenStudy (anonymous):

What would the answer be?

Nnesha (nnesha):

let's see multiply 4x and -2 by -3x \[\huge\rm \color{red}{-3x}(4x-2)\] \[\large \rm -3x \times 4x = ??\]

Nnesha (nnesha):

when you multiply same variables you should add their exponent \[\huge\rm x^m \times x^ n =x^{m+n}\]

Nnesha (nnesha):

-3x times 4x = ?? :-)